OpenStudy (anonymous):

Find the area of a triangle with a base of 8 m and a height of 6 m. A. 7 m2 B. 48 m2 C. 24 m2 D. 14 m2

OpenStudy (shamim):

c

OpenStudy (shamim):

can i explain

OpenStudy (anonymous):

yeah

OpenStudy (shamim):

ok

OpenStudy (shamim):

u know the area of a triange is= \[\frac{ 1 }{ 2 } \times base \times height\]

OpenStudy (shamim):

=\[\frac{ 1 }{ 2 } \times 8m \times 6m\]

OpenStudy (shamim):

=\[24m ^{2}\]

OpenStudy (anonymous):

ok thanks alot

OpenStudy (shamim):

welcome